# Controlled bit-flip of period-doubling and discrete time crystalline states in open systems

### v1

@@ Appendix A Equations of motion of the ODM
$da_{\mathrm{R}}=-\left(\omega a_{\mathrm{R}}+\frac{2\lambda}{\sqrt{N}}S^{x}+\kappa a_{\mathrm{I}}\right)+\sqrt{\frac{\kappa}{2}}dW_{2}$ (24b)
@@
where we expand the cavity mode in terms of its real and imaginary component, $a=a_{\mathrm{R}}+ia_{\mathrm{I}}$. In this form, the quantum fluctuations are captured by the two independent Wiener processes $W_{1}$ and $W_{2}$. They both satisfy the conditions $dW_{i}\right>=0$ and $dW_{i}dW_{j}\right>=\delta_{i,j}dt$ for $i=1,2$. For the DTWA, we instead consider the stochastic differential equations associated with each of the individual spins, $\vec{\sigma}_{i}$, appearing in Eq. (24) by writing $S^{x,y,z}=\sum_{i}^{N}\sigma_{i}^{x,y,z}$. This leads to the following equations of motion for the DTWA,
@@
$da_{\mathrm{R}}=-\left(\omega a_{\mathrm{R}}+\frac{2\lambda}{\sqrt{N}}\sum_{i}^{N}\sigma_{i}^{x}+\kappa a_{\mathrm{I}}\right)+\sqrt{\frac{\kappa}{2}}dW_{2}$ (25b)
@@
Appendix B Regimes with stable PD and DTC states
B.1 Resonance conditions for the parametric pendulum
@@
To identify the regimes in the DPP with PD states for periodic driving at $\omega_{d}=2\Omega$, we obtain the maximum value of $x=\sin(u)$ in the long-time limit. We do this by first initializing the system according to Eq. (16). We then allow the DPP to relax into a steady oscillating state within the time interval $t\in\left[0,200T_{d}\right]$. Finally, we obtain $\max[x]$ during the time interval $t\in\left[100T_{d},200T_{d}\right]$. If $\max[x]\geq x_{0}\approx 10^{-4}$, we consider the system to be in a PD state, otherwise the DPP only has a decaying oscillation towards $x=0$.
We present in Fig. 10 the $\max[x]$ of the DPP as a function of the driving amplitude $A$ for different dissipation strength $\gamma$. We can observe that $\max[x]$ has two different scaling behavior depending on whether it is on a parametric resonance or not. In particular, $\max[x]$ has a larger scaling exponent when it is in a PD state, which is most apparent for $\gamma=0.1\Omega$. The critical point separating these two regimes collapses when we consider the particular form of scaling for the driving amplitude
@@
With this scaling, the critical point lines up at $A_{\mathrm{scaled}}=1$. This suggests that when the DPP is resonantly driven at $\omega_{d}=2\Omega$, the minimum driving amplitude needed to push the system into a PD state is given to be
@@
B.2 Stability of PD and DTC states in the presence of noise
Figure 11: (a-b) Crystalline fraction of (a) the DPP and (b) the ODM for $\lambda_{0}<\lambda_{c}$ as a function of $\tilde{T}$ and $N$ respectively. The parameters considered for DPP are $\left\{\gamma,\omega_{d},\delta A\right\}=\left\{0.1\Omega,2\Omega,0.4\right\}$, while we consider $\left\{\kappa,\omega_{d},\delta A\right\}=\left\{1.0\omega,2\omega_{-},0.1\right\}$ for the ODM. The vertical dashed lines mark the values of $\tilde{T}$ and $N$ considered in the main text.
Noise can diminish the coherence of the PD and DTC states depending on its strength. This can either lead to thermally or quantum-activated switching between states [78] or the oscillations being washed out by strong fluctuations. To identify the regimes in which we can still test the robustness of bit-flip operations against noise without the need to account for these adverse effects, we quantify the quality of the oscillations associated with the PD and DTC states using the crystalline fraction $\chi$ defined as
@@ B.2 Stability of PD and DTC states in the presence of noise
where $P(\omega)$ is the power spectrum of the order parameter considered. Here, the relevant order parameters are the $x$-quadrature for the thermal DPP, and $S^{x}$ for the ODM.
@@
We present in Fig. 11 the crystalline fraction of the thermal DPP and ODM as a function of $\tilde{T}$ and $N$, respectively. We find that for small $\tilde{T}$ and large $N$, the thermal DPP and the ODM have a crystalline fraction of $\chi\approx 1$, indicating that the PD and DTC states at those regimes do not switch randomly and uncontrollably to their symmetry broken pairs due to strong noise. For $N=1/\tilde{T}\leq 10^{3}$, the crystalline fraction decreases, which means that the bit-flip operations cannot be tested in these values of $N$ and $\tilde{T}$ as they coincide with the noise-activated switching of PD and DTC states. As such, in the main text, we only considered noise strengths marked by the gray dashed lines, all of which are in the regime where $\chi\approx 1$. Note that in the case of the DTC states, the crystalline fraction do not approach $\chi=1$ in the thermodynamic limit due to the presence of higher harmonics that are integer multiples of its dominant frequency, $\omega_{d}/2$ [62].
Appendix C Defect protocol with phase error
@@
Figure 12: Absolute-time phase corresponding to the two degenerate DTC states of the ODM as a function of the phase of the drive $\theta$. The driving parameter considered are $\left\{\kappa,\omega_{d},A\right\}=\left\{1.0\omega,0.8\omega,0.55\right\}$. Note that the labeling of the DTC states are arbitrary, hence for this particular example, we label the state with the smaller $\varphi$ as the 00-bit state.
@@
All of the defect protocols considered in the main text involve the phase of the drive going back to zero after the defect. We address here what happens when the driving protocol does not go back to its original state after the defect protocol and instead obtains a phase shift $\theta_{f}$. We begin by determining the relationship between the degenerate DTC states and the phase of the driving protocol, $\theta$, in the absence of any defect. As shown in Fig. 12, the absolute-time phase $\varphi$ of the two degenerate DTC states follow a linear trend with $\theta$, indicating that $\theta$ dictates the possible values of $\varphi$. This can be understood by viewing the driving protocol as a periodic potential, in which the oscillations of the DTC are pinned. As such, the $\varphi$ of the two degenerate DTC states will follow the phase of the driving protocol accordingly.
@@
Figure 13: (a-b) Sketch of the bit-flip operation when the driving protocol has a phase error of (a) $\theta_{f}=0$, and (b) $\theta_{f}\neq 0$ after the defect protocol. (c) (Top panel) Exemplary PD state switching when the driving protocol has a phase error $\theta_{d}=\pi/2$ only during the defect protocol. (Bottom panel) PD state switching when the driving protocol has a phase error of $\theta_{f}=\pi/2$ after the defect protocol. For both dynamics, the parameters are $\omega_{d}=0.8$, $A=0.55$, $T_{r}=T_{\delta}=2.0T_{d}$.(d) Half-winding number distribution for the ODM under the continuous quenched-frequency protocol for the same parameters as Fig. 8(b).
@@
where $\theta_{i}$, $\theta_{d}$, and $\theta_{f}$ are the phases of the drive before, during, and after the defect protocol, respectively. Consider for instance $T_{r}=T_{\delta}$ such that $\omega_{d}^{\prime}$ is only dictated by the defect duration. Without loss of generality, we will set $\theta_{i}=0$. Suppose we consider a suitable $T_{\delta}$ such that the DTC switches after the defect when $\theta_{D}=\theta_{f}=0$. Given that the $\varphi$ of the two degenerate DTC states depends on the phase of the drive, we can infer that at $t>T_{r}$, the system does not perfectly switch into one of the original DTC states when $\theta_{f}\neq 2\pi n$ with $n\in\mathbb{Z}$, as shown in Fig. 13(b). Instead, the system incurs an absolute-time phase of $\varphi=\pi+\theta_{f}$, resulting in an imperfect bit-flip relative to the original bit encoded in the PD or DTC state. This behavior also implies that $\theta_{d}$ does not affect the overall bit-flip dynamics of the system. We verify this in Fig. 13(c), where we consider the cases of $\{\theta_{D},\theta_{f}\}=\{\pi/2,0\}$ and $\{\theta_{D},\theta_{f}\}=\{0,\pi/2\}$, respectively. As we can observe, we obtain perfect switching even when $\theta_{d}\neq 0$ provided $\theta_{f}=0$. As soon as $\theta_{f}$ becomes nonzero, the DTC obtains a phase shift of $\varphi=\pi+\theta_{f}$, which leads to an imperfect switching of a PD or DTC state.
Given that the bit-flip operation of DTC states only depends on $\theta_{f}$, we now present in Fig. 13(d) the $P_{w}$ of the ODM for the case of $T_{r}\neq T_{\delta}$ and $\theta_{f}=2\pi T_{r}/T_{\delta}$, with $\theta_{D}=0$. Unlike the driving protocol described in Eq. (21), this choice of $\theta_{f}$ preserves the continuity of the drive throughout the defect protocol. We use the same parameters as in Fig. 8(b) to construct Fig. 13(d). In contrast to the behavior of $P_{w}$ in Fig. 8(b), we find in Fig. 13(d) that $P_{w}$ has a continuous dependence with $\theta_{f}$, with the most probable $w$ following the linear trend $w=-\theta_{f}/2\pi=-T_{r}/T_{\delta}$, suggesting the sensitivity of the bit-flip operation in the final phase of the drive. Note that these results should apply to the thermal DPP as well since the same principles apply to them, i.e. the dependence of the degenerate PD states on the driving phase.

### v2

@@ Appendix A Equations of motion of the ODM
$da_{\mathrm{I}}=-\left(\omega a_{\mathrm{R}}+\frac{2\lambda}{\sqrt{N}}S^{x}+\kappa a_{\mathrm{I}}\right)+\sqrt{\frac{\kappa}{2}}dW_{2}$ (22b)
@@
where we expand the cavity mode in terms of its real and imaginary component, $a=a_{\mathrm{R}}+ia_{\mathrm{I}}$. In this form, the quantum fluctuations are captured by the two independent Wiener processes $W_{1}$ and $W_{2}$. They both satisfy the conditions $dW_{i}\right>=0$ and $dW_{i}dW_{j}\right>=\delta_{i,j}dt$ for $i=1,2$. For the DTWA, we instead consider the stochastic differential equations associated with each of the individual spins, $\vec{\sigma}_{i}$, appearing in Eq. (22) by writing $S^{x,y,z}=\sum_{i}^{N}\sigma_{i}^{x,y,z}/2$. This leads to the following equations of motion for the DTWA:
@@
$da_{\mathrm{I}}=-\left(\omega a_{\mathrm{R}}+\frac{\lambda}{\sqrt{N}}\sum_{i}^{N}\sigma_{i}^{x}+\kappa a_{\mathrm{I}}\right)+\sqrt{\frac{\kappa}{2}}dW_{2}$ (23b)
@@
Appendix B Regimes with stable period-doubled and DTC states
B.1 Resonance conditions for the parametric oscillator
@@
To identify the regimes in the PO with period-doubled states for $\omega_{d}=2\Omega$, we obtain the maximum value of $x=\sin(u)$ in the long-time limit. We do this by first initializing the system according to Eq. (14). We then allow the PO to relax into a steady oscillating state within the time interval $t\in\left[0,200T_{d}\right]$. Finally, we obtain $\max[x]$ during the time interval $t\in\left[100T_{d},200T_{d}\right]$. If $\max[x]\geq x_{0}\approx 10^{-4}$, we consider the system to be in a period-doubled state, otherwise, the PO only has a decaying oscillation towards $x=0$.
We present in Fig. 9 the $\max[x]$ of the PO as a function of the driving amplitude $A$ for different dissipation strength $\gamma$. We can observe that $\max[x]$ has two different scaling behaviors depending on whether it is on a parametric resonance or not. In particular, $\max[x]$ has a larger scaling exponent when it is in a period-doubled state, which is most apparent for $\gamma=0.1\Omega$. The critical point separating these two regimes collapses when we consider the particular form of scaling for the driving amplitude
@@
With this scaling, the critical point lines up at $A_{\mathrm{scaled}}=1$. This suggests that when the PO is resonantly driven at $\omega_{d}=2\Omega$, the minimum driving amplitude needed to push the system into a period-doubled state is given to be
@@
B.2 Stability of period-doubled and DTC states in the presence of noise
Figure 10: (a), (b) Crystalline fraction of (a) the PO and (b) the ODM for $\lambda_{0}<\lambda_{c}$ as a function of $\tilde{T}$ and $N$ respectively. The parameters considered for PO are $\left\{\gamma,\omega_{d},\delta A\right\}=\left\{0.1\Omega,2\Omega,0.4\right\}$, while we consider $\left\{\kappa,\omega_{d},\delta A\right\}=\left\{1.0\omega,2\omega_{-},0.1\right\}$ for the ODM. The vertical dashed lines mark the values of $\tilde{T}$ and $N$ considered in the main text.
Noise can diminish the coherence of the period-doubled and DTC states depending on its strength. This can either lead to thermally or quantum-activated switching between states [68] or the oscillations being washed out by strong fluctuations. To identify the regimes in which we can still test the robustness of bit-flip operations against noise without the need to account for these adverse effects, we quantify the quality of the oscillations associated with the period-doubled and DTC states using the crystalline fraction $\chi$ defined as
@@ B.2 Stability of period-doubled and DTC states in the presence of noise
where $P(\omega)$ is the power spectrum of the order parameter considered. Here, the relevant order parameters are the $x$-quadrature for the thermal PO, and $S^{x}$ for the ODM.
@@
We present in Fig. 10 the crystalline fraction of the thermal PO and ODM as a function of $\tilde{T}$ and $N$, respectively. We find that for small $\tilde{T}$ and large $N$, the thermal PO and the ODM have a crystalline fraction of $\chi\approx 1$, indicating that the period-doubled and DTC states at those regimes do not switch randomly and uncontrollably to their symmetry broken pairs due to strong noise. For $N=1/\tilde{T}\leq 10^{3}$, the crystalline fraction decreases, which means that the bit-flip operations cannot be tested in these values of $N$ and $\tilde{T}$ as they coincide with the noise-activated switching of the period-doubled and DTC states. As such, in the main text, we only considered noise strengths marked by the gray dashed lines, all of which are in the regime where $\chi\approx 1$. Note that in the case of the DTC states, the crystalline fraction does not approach $\chi=1$ in the thermodynamic limit due to the presence of higher harmonics that are integer multiples of its dominant frequency, $\omega_{d}/2$ [53].
Appendix C Fractal structures in bit-flip diagrams for weak dissipation
where $\vec{s}_{\mathrm{O},\mathrm{P}}=\begin{pmatrix}S^{x}_{\mathrm{O},\mathrm{P}}&S^{y}_{\mathrm{O},\mathrm{P}}&S^{z}_{\mathrm{O,P}}\end{pmatrix}^{T}/N$, and by the conservation of spin angular momentum for the ODM [62], $||\vec{s}_{O}||^{2}=||\vec{s}_{P}||^{2}=\frac{1}{4}$. The decorrelator determines the deviation between two identical systems, labeled as systems $\mathrm{O}$ and $\mathrm{P}$, initialized in a nearly identical state [30]. In particular, for regular dynamics, we expect that $\vec{s}_{\mathrm{O}}$ and $\vec{s}_{\mathrm{P}}$ will be aligned at all times and thus $d^{2}(t)=0$. Meanwhile, we get the maximum $d^{2}=1$ when $\vec{s}_{\mathrm{O}}$ and $\vec{s}_{\mathrm{P}}$ are anti-aligned with one another.
Appendix D Defect protocol with phase error
@@
Figure 12: Absolute-time phase corresponding to the two degenerate DTC states of the ODM as a function of the phase of the drive $\theta$. The driving parameter considered is $\left\{\kappa,\omega_{d},A\right\}=\left\{1.0\omega,0.8\omega,0.55\right\}$. Note that the labeling of the DTC states is arbitrary, hence, for this particular example, we label the state with the smaller $\varphi$ as the 00-bit state.
@@
All of the defect protocols considered in the main text involve the phase of the drive going back to zero after the defect. We address here what happens when the driving protocol does not go back to its original state after the defect protocol and instead obtains a phase shift $\theta_{f}$. We begin by determining the relationship between the degenerate DTC states and the phase of the driving protocol, $\theta$, in the absence of any defect. As shown in Fig. 12, the absolute-time phase $\varphi$ of the two degenerate DTC states follows a linear trend with $\theta$, indicating that $\theta$ dictates the possible values of $\varphi$. This can be understood by viewing the driving protocol as a periodic potential, in which the oscillations of the DTC are pinned. As such, the $\varphi$ of the two degenerate DTC states will follow the phase of the driving protocol accordingly.
@@
Figure 13: (a), (b) Sketch of the bit-flip operation when the driving protocol has a phase error of (a) $\theta_{f}=0$, and (b) $\theta_{f}\neq 0$ after the defect protocol. (c) (Top panel) Exemplary period-doubled state switching when the driving protocol has a phase error $\theta_{d}=\pi/2$ only during the defect protocol. (Bottom panel) Bit-flip when the driving protocol has a phase error of $\theta_{f}=\pi/2$ after the defect protocol. For both dynamics, the parameters are $\omega_{d}=0.8$, $A=0.55$, $T_{r}=T_{\delta}=2.0T_{d}$.(d) Half-winding number distribution for the ODM under the continuous quenched-frequency protocol for the same parameters as Fig. 7(b).
@@
where $\theta_{i}$, $\theta_{d}$, and $\theta_{f}$ are the phases of the drive before, during, and after the defect protocol, respectively. Consider for instance $T_{r}=T_{\delta}$ such that $\omega_{d}^{\prime}$ is only dictated by the defect duration. Without loss of generality, we will set $\theta_{i}=0$. Suppose we consider a suitable $T_{\delta}$ such that the DTC switches after the defect when $\theta_{D}=\theta_{f}=0$. Given that the $\varphi$ of the two degenerate DTC states depends on the phase of the drive, we can infer that at $t>T_{r}$, the system does not perfectly switch into one of the original DTC states when $\theta_{f}\neq 2\pi n$ with $n\in\mathbb{Z}$, as shown in Fig. 13(b). Instead, the system incurs an absolute-time phase of $\varphi=\pi+\theta_{f}$, resulting in an imperfect bit-flip relative to the original bit encoded in the period-doubled or DTC state. This behavior also implies that $\theta_{d}$ does not affect the overall bit-flip dynamics of the system. We verify this in Fig. 13(c), where we consider the cases of $\{\theta_{D},\theta_{f}\}=\{\pi/2,0\}$ and $\{\theta_{D},\theta_{f}\}=\{0,\pi/2\}$, respectively. As we can observe, we obtain perfect switching even when $\theta_{d}\neq 0$ provided $\theta_{f}=0$. As soon as $\theta_{f}$ becomes nonzero, the DTC obtains a phase shift of $\varphi=\pi+\theta_{f}$, which leads to an imperfect switching of a period-doubled or DTC state.
Given that the bit-flip operation of DTC states only depends on $\theta_{f}$, we now present in Fig. 13(d) the $P_{w}$ of the ODM for the case of $T_{r}\neq T_{\delta}$ and $\theta_{f}=2\pi T_{r}/T_{\delta}$, with $\theta_{D}=0$. Unlike the driving protocol described in Eq. (19), this choice of $\theta_{f}$ preserves the continuity of the drive throughout the defect protocol. We use the same parameters as in Fig. 7(b) to construct Fig. 13(d). In contrast to the behavior of $P_{w}$ in Fig. 7(b), we find in Fig. 13(d) that $P_{w}$ has a continuous dependence with $\theta_{f}$, with the most probable $w$ following the linear trend $w=-\theta_{f}/2\pi=-T_{r}/T_{\delta}$, suggesting the sensitivity of the bit-flip operation in the final phase of the drive. Note that these results should apply to the thermal PO as well since the same principles apply to them, i.e. the dependence of the degenerate period-doubled states on the driving phase.